OpenStudy (lgbasallote):

LGBADERIVATIVE: \[\huge y = \cos^2 x\] y " = ?

OpenStudy (lgbasallote):

\[y' = -2\sin x \cos x\]

OpenStudy (lgbasallote):

\[y'' = 2\sin ^2 x - 2 \cos ^2 x?\]

OpenStudy (lgbasallote):

\[y'' = 2(\cos (2x) )?\]

OpenStudy (zarkon):

close

OpenStudy (lgbasallote):

so it's wrong?

OpenStudy (zarkon):

yes...just a little

OpenStudy (callisto):

\[cos^2 x = \frac{cos2x + 1}{2}\] \[y' = \frac{d}{dx}cos^2 x = \frac{d}{dx}\frac{cos2x + 1}{2} = \frac{1}{2}(-sin2x)(2) = -sin2x\] \[y'' = \frac{d}{dx} -sin2x = -2cos2x\] I haven't done diff. for long ... Sorry :(

OpenStudy (lgbasallote):

wait....i think it should be 2... @Zarkon ?

OpenStudy (lgbasallote):

\[2\sin^2 x + 2\cos ^2 x\] \[2(1)\]

OpenStudy (zarkon):

no

OpenStudy (lgbasallote):

callisto and i had same answer :/

OpenStudy (zarkon):

nope

OpenStudy (callisto):

Not the same :|

OpenStudy (zarkon):

close, but not the same

OpenStudy (lgbasallote):

huh? what's the difference between \[-2 (\cos (2x) )\] and \[-2 \cos 2x\] ???

OpenStudy (zarkon):

you didn't put a - in your answer above

OpenStudy (callisto):

You didn't type the negative sign.

OpenStudy (lgbasallote):

oops lol

OpenStudy (lgbasallote):

it shouldve been \[-2\sin^2 x + 2\cos^2 x?\]

OpenStudy (anonymous):

what is a \(-\) between friends?

OpenStudy (lgbasallote):

hmm wonder what